1. Brockman Corporation's earnings per share were $3.50 last year, and its growth rate during the prior 5 years was 7.8% per year. If that growth rate were maintained, how many years would it take for Brockman's EPS to triple?

Select the correct answer.

a. 16.23
b. 14.63
c. 17.83
d. 19.43
e. 13.03

2. A new investment opportunity for you is an annuity that pays $3,900 at the beginning of each year for 3 years. You could earn 5.5% on your money in other investments with equal risk. What is the most you should pay for the annuity?

Select the correct answer.

a. $11,138.85
b. $11,100.65
c. $11,177.05
d. $11,157.95
e. $11,119.75
Answer #1

1.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

(3*3.5)=3.5*(1.078)^n

3=(1.078)^n

Taking log on both sides;

log 3=n*log (1.078)

n=log 3/log (1.078)

=14.63 years(Approx).

2.

Present value of annuity due=(1+rate)*Annuity[1-(1+interest rate)^-time period]/rate

=1.055*3900[1-(1.055)^-3]/0.055

=3900*2.84631972

=$11100.65(Approx).